Already famous by the time he arrived at Princeton University in 1933, Einstein had suggested a salary of $3,000. The university didn’t want other scholars being paid more than Einstein and increased the salary to $10,000. The CPI in 1933 is 13 and the CPI in 2017 is 245. What’s Albert’s salary in 2017?
Answer #1

Albert’s salary in 2017 = Salary in 1933*(CPI in 2017/CPI in 1933) = (10,000)*(245/13) = 2,450,000/13 = 188,461.54

So, Albert's salary in 2017 is $188,461.54